OpenStudy (anonymous):

Write an equation of a line that passes through the point (2, 3) and is parallel to the line y = 2/3 x + 5.

OpenStudy (anonymous):

y = -2/3 x y =-2/3 x − 7 y = 3/2 x y = 3/2 x + 7

OpenStudy (geekfromthefutur):

My sis told me u need help

OpenStudy (anonymous):

yeah could you help me this question

OpenStudy (anonymous):

y = 2 3x + 5 Is this the equation of the line?

OpenStudy (anonymous):

i do not know

OpenStudy (anonymous):

What is the first equation the question gives you : ) Sorry, it is a bit hard to interpret your question ><

OpenStudy (anonymous):

Write an equation of a line that passes through the point (2, 3) and is parallel to the line y = 2/3 x + 5. this is the question so i think the first equation would be y=2/3x+5

OpenStudy (anonymous):

Okay, so your previous working out, involves making y the subject, is correct notice how the coefficient in front of x is the gradient

OpenStudy (anonymous):

So from: y = 3/2 x + 7 m = 3 /2 With this passing through the point (2,3) remember your point-gradient formula y - y1 = m (x -x1)

OpenStudy (anonymous):

what is y and what is x?

OpenStudy (anonymous):

wait is the answer y+3/2 x + 7?

OpenStudy (anonymous):

x1 and y1 are from your point (2,1) where 2 represents the x value and 1 is the y value

OpenStudy (anonymous):

I haven't worked it out, give me a moment

rvc (rvc):

to revise : Basic Straight Lines http://openstudy.com/study#/updates/554cca3ee4b075612122616a go to the above link :)

OpenStudy (anonymous):

I got y = 3x - 4 2

OpenStudy (anonymous):

A: y = -2/3 x B:y =-2/3 x − 7 C:y = 3/2 x D:y = 3/2 x + 7 these are the answer choices

OpenStudy (anonymous):

I think we may have misinterpreted out gradient value of 3/2, I rechecked it again an i got 2/3 through working out, my outcome was: y = 2x -1 3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Alright no worries